A pole that is 3.5m tall casts a shadow that is 1.09m long. At the same time, a nearby tower casts a shadow that is 45.25m long. How tall is the tower? Round your answer to the nearest meter.


Answer by math_tutor2020(3817)


Answer: 145 meters

Work Shown:
(pole height)/(pole's shadow) = (tower's height)/(tower's shadow)
(3.5 meters)/(1.09 meters) = (x meters)/(45.25 meters)
(3.5)/(1.09) = x/(45.25)
x/(45.25) = (3.5)/(1.09)
x = (45.25)*(3.5)/(1.09)
x = 145.298165137614 approximately
x = 145 meters